Grade 12th passGeneral Physics

A shot after leaving the gun at around passes over a wall horizontally . if the wall is 10 m high and 30 m away from the gun find the velocity of projection and direction of shot

sayali gavali
8 Years agoGrade 12th pass

1 Answer

Eshan
8 Years ago
Dear student,

Consider the motion in two coordinate axes, horizontal and vertical ones.

In the vertical direction,

0^2-u_y^2=2(-10)(10)
\implies u_y=10\sqrt{2}m/s

Time in which the particle crosse the wall=\dfrac{u_y}{g}=\sqrt{2}s
Hence the horizontal component of the velocity of projection isu_x=\dfrac{30}{\sqrt{2}}m/s=15\sqrt{2}m/s
Hencetan\theta=\dfrac{u_y}{u_x}=\dfrac{2}{3}

\implies u=5\sqrt{26}m/s